OpenStudy (anonymous):

What is the vertex of the graph of y = −2x^2 + x + 3?

OpenStudy (anonymous):

general form of this type of equation:\[y=ax ^{2}+bx +c\] vertex of this parabola is: -b/2a in your case a=-2, b=1

OpenStudy (kainui):

The derivative of a graph shows you the slope at any point on the other graph. So where the slope of the graph is 0, then you have the vertex.

OpenStudy (kainui):

^ assuming you're in calculus.

OpenStudy (anonymous):

:)

OpenStudy (saifoo.khan):

Method 3: covert the equation into vertex form. ;)

OpenStudy (anonymous):

-1/4

OpenStudy (anonymous):

?

OpenStudy (anonymous):

i need another fraction though -1/4 isnt the only one how do i get the other fraction

OpenStudy (anonymous):

@myko

OpenStudy (anonymous):

@Kainui

OpenStudy (anonymous):

the answer should be 1/4, not -1/4. And the parabola has only one vertex

OpenStudy (anonymous):

but the have it as (1/4,25/8) or (1/4,27/8). which one do i choose @myko

OpenStudy (anonymous):

the choices i meant

OpenStudy (anonymous):

it is (x,y) so find y value for x=1/4. For that plug x=1/4 in to the equation

OpenStudy (anonymous):

ok i can do that thanks

OpenStudy (anonymous):

y = −2*(1/4)^2 + (1/4) + 3

OpenStudy (anonymous):

25/8 is th y value!!!! thanks myko

OpenStudy (anonymous):

yw